OpenStudy (anonymous):

please help me. ACTUALLY. What is the derivative of: f(t)=20[1-cos^29t/1000pi)] PLEEEASE

OpenStudy (anonymous):

write it correctly first. theres no such thing as cos^29t

OpenStudy (anonymous):

cos^2(t/1000pi) i guess i didnt hit shift when I hit 9.

OpenStudy (anonymous):

\[f(t) = 20 - 20 \cos(\frac{t}{1000\pi})\] the derivative of cos(x) is -sin(x), the derivative of 20 is 0, so \[f'(t) = 0 - (-20\sin(\frac{t}{1000\pi}) \frac{1}{1000\pi})\] where 1/1000pi is the derivative of t/1000 pi (chain rule!). So the result is \[f'(t) =\frac{1}{50\pi}\sin(\frac{t}{1000\pi}) \]